# Shaping Sparse Rewards in Reinforcement Learning: A Semi-supervised Approach

### v3

@@ Introduction
Parameter Confidence threshold $\lambda$, loss weight $\alpha$, set $Z$ with size $N_{z}$, episodes $T$, update probability $p_{u}$;
@@
5:   if $r_{t}\notin\{r_{1},\dots,r_{t-1}\}$ then
6:    Obtain $N_{z}$ new $z_{i}$ by interpolating from all historical observation $r_{1},...,r_{t}$
@@
9:    Calculate confidence $\mathbf{q_{t}}(\theta)$, $\mathbf{q_{t}^{w,\,\prime}(\theta)}$ of non-zero and zero reward transitions over set $Z$ under parameter $\theta$.
10:    Perform reward shaping on zero-reward transitions in $\mathcal{D}$ of the $p_{u}$ ratio.
@@
12:   Choose a batch of transitions $\mathcal{B}$ from $\mathcal{D}$ and update $Q$ function with shaped reward $\hat{r}$,
$Q(s,a)\leftarrow Q(s,a)+\alpha\cdot\frac{1}{|\mathcal{B}|}\\ \cdot\sum_{(s,a,\hat{r},s^{\prime})\in\mathcal{B}}\left(\hat{r}+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\right).$
@@
14:  Update parameter $\theta$ through minimizing the objective $L=L_{QV}+\alpha L_{s}+(1-\alpha)L_{r}$
@@
Experiment Results
(a) Montezuma
Revenge
(b) Pitfall
(c) Fetch and Reach
Figure 2: (a) and (b) show the game scenes from Montezuma Revenge and Pitfall as Sparse reward environment of the Arcade Learning Environment (ALE), commonly referred to as Atari. (c) is a snapshot of robotic arm performing Fetch and Reach task in Gymnasium-Robotics environment.
(a) Seaquest
(b) Hero
@@
Figure 3: Figures (a–d) demonstrate the best score in 1000 epochs of different algorithms in 4 Atari environments (mean $\pm$ std over 3 seeds). Different curves represent the SSRS framework and its variants, as well as RCP and SAC algorithms. The legends of each sub figure are uniformly labeled in Figure(a)
@@
The experiments aim to validate whether the SSRS using semi-supervised learning is more efficient in reward shaping than the self-supervised method, thereby improving the best score achieved by the agent in both the Atari game environment (Bellemare et al. 2013) and the robotics manipulation environment (Plappert et al. 2018) (see Figure 2). To demonstrate the superiority of double entropy data augmentation over other data augmentation methods in non-visual observation tasks, we use Random Access Memory (RAM) as the observation. Additionally, ablation experiments are conducted to demonstrate the importance of monotonicity constraints in the application of SSL techniques. We will analyze the characteristics of the trajectory space through experiments to provide insights into the underlying reasons of such improvement.
@@
Experimental Setup. RAM observation refers to a representation of the Atari RL environments’ internal state directly from its Random Access Memory, typically 128 bytes of RAM (i.e., a vector of 128 integers, each in the range [0, 255]). The implementation for this experiment is based on the Tianshou RL library (Weng et al. 2022). Across three experimental seeds, a maximum of 2000 transition samples per episode are collected and appended to the buffer $\mathcal{D}$, a buffer with a capacity of 30k transitions. The maximum test score achieved from the start of training up to the current epoch, along with its variance across different experimental seeds, is recorded and plotted in the figures throughout this section. Some of the hyperparameter settings are as Table 4 in Appendix, which are determined based on preliminary hyperparameter tuning results, and consistent hyperparameter settings are used in both the Robotics and Atari experiments.
Performance of SSRS
@@
Figure 4: This figure shows the best score curve of SSRS variants and aforementioned baseline over 500 epochs in the robotic manipulation environment FetchReach (mean $\pm$ std over 3 seeds).
@@
To evaluate the proposed SSRS framework, we compare it with several baseline approaches including supervised-based methods (RCP (Kumar, Peng, and Levine 2019) and SORS (Memarian et al. 2021)) as well as backbone reinforcement learning algorithms in both Atari and robotic environments. Since this paper focuses on reward shaping, the backbone algorithms used are not the main emphasis. The backbone algorithms for Atari and Robotic environments are Soft Actor-Critic (SAC, Haarnoja et al. (2019)) and Deep Deterministic Policy Gradient (DDPG, Lillicrap et al. (2015)), respectively. Several variants of the SSRS framework are employed, including SSRS-S, which uses double entropy augmentation, SSRS-C, which adopts the cutout method for data augmentation, and SSRS-M, which adopts the smooth method for data augmentation.
@@
In Atari environment, the SSRS algorithm and its variants outperform the baseline algorithm and supervised-based RCP across all four games (see Figure 3). This is because When the proportion of transitions with non-zero rewards is low (i.e., when $\mu$ is small), the self-supervised method has fewer samples containing ”ground truth labels” available for learning. In such cases, the SSRS approach, which leverages semi-supervised learning to infer trajectory decision boundaries, can estimate rewards from transitions more effectively than the self-supervised method. (The concept of trajectory decision boundaries will be discussed in section ’Feasibility of Semi-Supervised Learning’.) Seaquest and Hero games are considered to be moderately reward-sparse, so we add them to the setting to evaluate the general performance of SSRS. In terms of convergence speed, the SSRS-S variants outperform RCP in the Venture environment. Moreover, in games with extremely sparse rewards, such as MonteZumaRevenge, SSRS significantly outperforms the RCP algorithm, achieving nearly 2 times the best score in MonteZumaRevenge (see Figure 3(c)). In the FetchReach environment in Robotics, where the reward signal is binary, i.e., the reward is -1 if the end effector hasn’t reached its final target position, and 0 if the end effector is in the final target position, the SSRS framework outperforms the self-supervised method SORS (see Figure 4).
@@
We also conducte experiments to evaluate the performance of SSRS when using different data augmentation methods as perturbations for SSL, aiming to investigate whether the proposed double entropy data augmentation provides greater improvement in terms of final best score for SSRS compared to other traditional data augmentation methods (e.g., flip, scale, translate). The detailed results and analysis are provided in Appendix.
@@
Figure 5: (a) represents SSRS with a monotonicity constraint, (b) represents SSRS without a monotonicity constraint, and (c) represents the baseline algorithm DQN. Shaped rewards from the buffer $\mathcal{D}$ were sampled at epochs 200, 400, 600, 800, and 1000. Since the reward value distribution is sparse, the logarithm of the reward values is used as the x-axis, the y-axis represents values of epoch scale to 1/200, and the z-axis represents the normalized reward distribution probability.
Ablation Experiments of Monotonicity Constraint
We conduct a series of experiments on the SSRS framework under conditions with and without monotonicity constraint and study the distribution of reward values at different stages of training in both scenarios.
The average best score obtained is presented in Table 1. In three environments, SSRS with a monotonicity constraint achieves higher best scores compared to SSRS without a monotonicity constraint. We find that adding the monotonicity constraint reduces the variance as well as the final best score in reward-sparse environments with long-term goals (see Figure 3 and 1 in Appendix). One of the reasons of such difference in sparse and less-sparse scenarios is our strategy of configure in update probability $p_{u}$. The hyperparameter $p_{u}$ is set proportional to the number of non-zero-reward transitions, $\mathcal{N}_{r}$. However, the variation in the number of non-zero-reward transitions across different environments is not linear. Although we improved the method by setting $p_{u}$ to the order of $\log(\mathcal{N}_{r})$ during the early stages of training, this adjustment does not cover all possible scenarios.
Further investigation of the reward distribution in Hero (see Figure 5) reveals that, according to the reward distribution of the backbone algorithm without reward shaping, most of the reward signals in this environment consist of small reward values, with a lack of medium to long-term reward feedback. SSRS without a monotonicity constraint only provides reward estimates close to the mean. However, this deviates from the actual reward distribution of the environment, causing the reward estimator to degrade into a binary classifier. Experimental results also indicate a performance drop without the constraint. In contrast, SSRS with a monotonicity constraint shapes a reward distribution that closely approximates the true reward distribution while also exhibiting the properties of a positively skewed normal distribution.
Feasibility of Semi-Supervised Learning
@@
Figure 6: This figure shows the consensus matrix of reward distribution in Hero after 1000 episodes iterations, which indicates that the distribution of trajectories along the reward dimension in the trajectory space exhibits a certain clustering property. The cluster method used is Gaussian Mixture Models, and the number of iterations for consesus matrix is 100.
@@
To further investigate why SSL method can infer rewards from trajectories more effectively than self-supervised methods—and why SSL techniques can be applied to reward shaping, we conduct the following experiment and analyses. We examine the trajectory distribution and obtain the consensus matrix $\mathcal{C}$ in Figure 6. Each element $a_{ij}$ $\in\mathcal{C},0\leq a_{ij}\leq 1,$ is frequency at which the two trajectories $\tau_{i}$ and $\tau_{j}$ are assigned to the same shaped reward value. Note that we use estimated reward values for labels.
@@
Larger estimated reward values, i.e., rewards given for achieving long-term goals, exhibit clearer diagonal boundaries compared to smaller reward values. However, the decision boundaries inside high reward values are more ambiguous, with only one clear diagonal block, aligning with the clustering assumption of SSL. SSL methods relies on the smoothness assumption and clustering assumption (Yang et al. 2023) on the augmented data, in this case, the trajectories. To be specific, for classification models, the smoothness assumption indicates if two points $x_{1},x_{2}$ reside in a high-density region are close, then so should be their corresponding outputs $y_{1},y_{2}$. And clustering assumption indicates that samples from the same class are closer to each other, while there are significant boundaries between samples from different classes and the decision boundary of the model should be as far as possible from regions of high data density. Notably, this observed ambiguity in high reward values has minimal practical impact on SSRS, as the reward estimator outputs these high rewards with negligible confidence probabilities.
@@
Unlike conventional data augmentation techniques that may introduce excessive perturbations to non-image data—violating the smoothness and clustering assumptions— double entropy augmentation better preserves these assumptions in trajectory space. This is crucial to the step of calculating estimated reward, where the $V$ network and $Q$ network are not able to separate trajectories to learn a good representation of the space. From the perspective of the teacher-student model (Tarvainen and Valpola 2017), such a ‘muddled’ teacher would lead the student, i.e. policy module, to learn an unstable policy.
@@
By also referring to Figure 5, it can be observed that the three ‘phases’ of significance in the consensus matrix plot correspond to the downward trend of the reward distribution probability. This can be understood from the perspective that as the estimated reward value increases, the randomness of the trajectory decreases, leading to greater clustering significance in the trajectory space, and vice versa.
Conclusion
@@
In this paper, we propose a Semi-Supervised Reward Shaping (SSRS) framework which utilizes zero reward trajectories by employing SSL technique. Additionally, we propose the double entropy data augmentation method for consistency regularization and a monotonicity constraint over modules of reward estimator. Our model outperforms previous supervised-based RCP and SORS methods in Atari environments, with a maximum of twice increase in reaching higher best scores. Moreover, the double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods. With SSL techniques, we can deploy agents on significantly sparser trajectory data. From a reverse perspective, it is possible to obtain learnable trajectories by artificially annotating rewards for a minimal number of transitions within a large set of trajectory data.

### v4

@@ 3.4 Entropy Augmentation for SSRS
Parameter Confidence threshold $\lambda$, reward set cardinality $N_{z}$, episodes $T$, consistency coefficient $\beta$, update probability $p_{u}$;
@@
5:   if $r_{t}\notin\{0,r_{1},\dots,r_{t-1}\}$ and $|Z|<N_{z}$ then
6:    Append $r_{t}$ to reward set $Z\leftarrow Z\cup\{r_{t}\}$
@@
9:    Calculate confidence $\mathbf{q_{t}}(\theta)$, $\mathbf{q_{t}^{w,\,\prime}(\theta)}$ of non-zero and zero reward transitions over set $Z$ under parameter $\theta$
10:    Perform reward shaping on zero-reward transitions in $\mathcal{D}$ of the $p_{u}$ ratio
@@
12:   Choose a batch of transitions $\mathcal{B}$ from $\mathcal{D}$ and update $Q$ function with shaped reward $\hat{r}$, and set step-size $\eta$:
$Q(s,a)\leftarrow Q(s,a)+\frac{\eta}{|\mathcal{B}|}\\ \sum_{(s,a,\hat{r},s^{\prime})\in\mathcal{B}}\bigl(\hat{r}+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\bigr).$
@@
14:  Update parameter $\theta$ through minimizing the objective $L=L_{QV}+\beta L_{s}+(1-\beta)L_{r}$
@@
4 EXPERIMENT RESULTS
(a) Kangaroo
(b) Seaquest
@@ 4 EXPERIMENT RESULTS
(e) Hero
Figure 2: (a)-(b) Reward curves of SSRS using C51 as the backbone, compared with the RCP algorithm. (c)-(e) Best score curves of SSRS under different data augmentation methods (using SAC as the backbone), compared with RCP. Note: To avoid cluttered curves from multiple variants, the best score—defined as the maximum score achieved from the beginning up to the current test episode—is plotted in (c)-(e). All curves represent the mean ± standard deviation over 5 seeds. The horizontal axis denotes test steps, and training is conducted for 1000k frames.
@@
The experiments aim to validate whether the SSRS using semi-supervised learning is more efficient in reward shaping than the self-supervised method, thereby improving the best score achieved by the agent in both the Atari game environment Bellemare et al. (2013) and the robotic manipulation environment Plappert et al. (2018). To demonstrate the superiority of entropy augmentation over other data augmentation methods in non-visual observation tasks, we use Random Access Memory (RAM) as the observation. Additionally, ablation experiments are conducted to demonstrate the importance of monotonicity constraints. We will analyze the characteristics of the trajectory space through experiments to provide insights into the underlying reasons of such improvement.
@@
Experimental Setup. RAM observation refers to a representation of the Atari RL environments’ internal state directly from its Random Access Memory, typically 128 bytes of RAM (i.e., a vector of 128 integers, each in the range [0, 255]). The implementation for this experiment is based on the Tianshou RL library Weng et al. (2022). Across experimental seeds, a maximum of 10000 transition samples per episode are collected and appended to the buffer $\mathcal{D}$, a buffer with a capacity of 30k transitions. Some of the hyperparameter settings are as Table. 4 in Appendix, which are determined based on preliminary hyperparameter tuning results, and consistent hyperparameter settings are used in both the Robotics and Atari experiments.
4.1 Performance of SSRS
@@
Figure 3: This figure shows the best score curve of SSRS variants and aforementioned baseline in the robotic manipulation environment FetchReach (mean $\pm$ std over 3 seeds). The horizontal axis denotes test steps, and training is conducted for 1000k frames.
@@
To evaluate the proposed SSRS framework, we compare it with several baseline approaches including supervised-based methods (RCP Kumar et al. (2019) and SORS Memarian et al. (2021)) as well as backbone reinforcement learning algorithms in both Atari and Robotics environments. Since this paper focuses on reward shaping, the backbone algorithms used are not the main emphasis. The backbone algorithms for Atari and Robotics environments are Categorical DQN (C51, Bellemare et al. (2017)), Soft Actor-Critic (SAC, Haarnoja et al. (2019)) and Deep Deterministic Policy Gradient (DDPG, Lillicrap et al. (2015)). Several variants of the SSRS framework are employed, including SSRS-S, which uses entropy augmentation, SSRS-C, which adopts the cutout method for data augmentation, and SSRS-M, which adopts the smooth method for data augmentation.
We selected five Atari games known for their sparse reward structures. For instance, in Seaquest, the player controls a submarine to rescue divers while avoiding underwater threats, with meaningful rewards occurring only upon successful rescues or eliminations. Similarly, in Kangaroo, the kangaroo character must navigate platforms to rescue its offspring, with rewards primarily tied to reaching specific objectives. Since these Atari environments operate on discrete action spaces, we first compare our method with the RCP algorithm, which is designed for discrete action settings. Comparisons with the SORS algorithm, developed for continuous action spaces, are deferred to later experiments in Robotics environments. The resulting learning curves and numerical results are presented in the figure and table below.
@@
the SSRS algorithm and its variants outperform the baseline algorithm and supervised-based RCP across all five games (see Figure. 2). This is because When the proportion of transitions with non-zero rewards is low, the self-supervised method has fewer samples containing ”ground truth labels” available for learning. In such cases, the SSRS approach, which leverages semi-supervised learning to infer trajectory decision boundaries, can estimate rewards from transitions more effectively than the self-supervised method. (The concept of trajectory decision boundaries will be discussed in section ’Feasibility of Semi-Supervised Learning’.) Seaquest and Hero games are considered to be moderately reward-sparse, so we add them to the setting to evaluate the general performance of SSRS. In terms of convergence speed, the SSRS-S variants outperform RCP in the Venture environment. Moreover, in games with extremely sparse rewards, such as MonteZumaRevenge, SSRS significantly outperforms the RCP algorithm, achieving nearly 2 times the best score in MonteZumaRevenge (see Figure. 2(c)). In the FetchReach environment in Robotics, where the reward signal is binary, i.e., the reward is -1 if the end effector hasn’t reached its final target position, and 0 if the end effector is in the final target position, the SSRS framework outperforms the self-supervised method SORS (see Figure. 3).
@@
We also examine the performance of SSRS when using different data augmentation methods as perturbations for SSL, aiming to investigate whether the proposed entropy augmentation provides greater improvement in terms of final best score for SSRS compared to other traditional data augmentation methods (e.g., flip, scale, translate). The detailed results and analysis are provided in Appendix. Additionally, we conduct a series of experiments on the SSRS framework under conditions with and without monotonicity constraint. The average best score obtained is presented in Table. 1. In three environments, SSRS with a monotonicity constraint achieves higher best scores compared to SSRS without a monotonicity constraint.
@@
4.2 Feasibility of Semi-Supervised Learning
@@
Figure 4: This figure shows the consensus matrix of reward distribution in Hero after 1000 episodes iterations, which indicates that the distribution of transitions along the reward dimension in the trajectory space exhibits a certain clustering property. The cluster method used is Gaussian Mixture Models, and the number of iterations for consensus matrix is 100.
@@
To further investigate why SSL method can infer rewards from transitions more effectively than self-supervised methods—and why SSL techniques can be applied to reward shaping, we conduct the following experiment and analysis. We examine the transition distribution and obtain the consensus matrix $\mathcal{C}$ in Figure. 4. Each element $a_{ij}$ $\in\mathcal{C},0\leq a_{ij}\leq 1,$ is frequency at which the two transitions $\tau_{i}$ and $\tau_{j}$ are assigned to the same shaped reward value. Note that we use estimated reward values for labels.
@@
Larger estimated reward values, i.e., rewards given for achieving long-term goals, exhibit clearer diagonal boundaries compared to smaller reward values. However, the decision boundaries inside high reward values are more ambiguous, with only one clear diagonal block, aligning with the clustering assumption of SSL. SSL methods relies on the smoothness assumption and clustering assumption Yang et al. (2023) on the augmented data, in this case, the trajectories. To be specific, for classification models, the smoothness assumption indicates if two points $x_{1},x_{2}$ reside in a high-density region are close, then so should be their corresponding outputs $y_{1},y_{2}$. And clustering assumption indicates that samples from the same class are closer to each other, while there are significant boundaries between samples from different classes and the decision boundary of the model should be as far as possible from regions of high data density. Notably, this observed ambiguity in high reward values has minimal practical impact on SSRS, as the reward estimator outputs these high rewards with negligible confidence probabilities.
@@
Unlike conventional data augmentation techniques that may introduce excessive perturbations to non-image data—violating the smoothness and clustering assumptions— entropy augmentation better preserves these assumptions in trajectory space. This is crucial to the step of calculating estimated reward, where the $R_{V}$ network and $R_{Q}$ network are not able to separate transitions to learn a good representation of the space. From the perspective of the teacher-student model Tarvainen and Valpola (2017), such a ‘muddled’ teacher would lead the student, i.e. policy module, to learn an unstable policy.
@@
From Figure. 4, it can be observed that the three ‘phases’ of significance in the consensus matrix plot correspond to the downward trend of the reward distribution probability. This can be understood from the perspective that as the estimated reward value increases, the randomness of the transition decreases, leading to greater clustering significance in the trajectory space, and vice versa.
4.3 Hyperparameter Robustness Study
SSRS introduces three hyperparameters, the update probability $p_{u}$, confidence threshold $\lambda$, and loss coefficient $\beta$. For supervised-based reward shaping Memarian et al. (2021); Kumar et al. (2019); Peng et al. (2019), the parameter controlling the update portion is natural, and experimental results (see Figure. 5) also demonstrate its limited impact on the robustness of this type of model, except when $p_{u}$ approaches zero. In this case, no transitions undergo reward shaping, thus making SSRS unattainable to improve performance. The confidence threshold $\lambda$ is one of the few hyperparameters in pseudo-labeling models Lee (2013); Sohn et al. (2020), and we also investigate its influence under the reinforcement learning framework. Similar to that observed in computer vision tasks, performance peaks occur when the confidence threshold value is between 0.8-0.9. Finally, the loss coefficient $\beta$ has been shown (see Table. 2) to primarily affect the training process, with minimal impact on the final results.
Figure 5: This figure illustrates the impact of the hyperparameters update probability $p_{u}$ and confidence threshold $\lambda$ on SSRS (using the Venture environment as an example). The lower x-axis represents the values of the confidence threshold, while the upper x-axis represents the values of the update probability.
Table 2: Performance of SSRS with different consistency coefficients $\beta$ in the Venture environment. This table shows the best scores achieved and the average corresponding training epochs when varying the values of $\beta$.
Consistency Coefficient $\beta$ 0.3 0.4 0.5 0.6 0.7
Best score 226 230 223 233 236
In epoch 854 853 862 877 892
5 Conclusion
@@
In this paper, we propose a Semi-Supervised Reward Shaping (SSRS) framework which utilizes zero reward trajectories by employing SSL technique. Additionally, we propose the entropy augmentation method for consistency regularization and a monotonicity constraint over modules of reward estimator. Our model outperforms previous supervised-based RCP and SORS methods in Atari environments, with a maximum of twice increase in reaching higher best scores. Moreover, the entropy augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods.